OpenStudy (anonymous):

How do I get the integral of e^(-x / 2)

OpenStudy (freckles):

u=-x/2 then du=-1/2dx

OpenStudy (solomonzelman):

he is right (obviously) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}e^{-x/2}~dx}\) , you are setting \(\large\color{slate}{u=-x/2}\) . If you want, I can show a couple of example of a u sub if you feel uncomfortable talking this integral.

OpenStudy (solomonzelman):

examples*